OpenStudy (anonymous):

Select the point that is on the right side of the vertex of y = |x + 4| + 3. (–3, 4) (–4, 3) (–5, 4) (–6, 5)

OpenStudy (anonymous):

When working with absolute values, remember that the quantity inside the | | has two values, one positive and one negative, so there will be two possible solutions. Set the quantity to each value and solve for the unknown. y = |x + 3| + 4 y = ± (x + 3) + 4 If y = (x + 3) + 4, y = x + 7 If y = - (x + 3) + 4, y = - x - 3 + 4 y = - x + 1 y = y, so x + 7 = - x + 1 x + x = 1 - 7 2x = - 6 x = - 6 / 2 x = - 3 and y = - 3 + 7 y = 4 Point of Intersection (- 3, 4)

OpenStudy (anonymous):

is it okay friend

OpenStudy (anonymous):

thanku